OpenStudy (thefluffmuncher):

2. Select the proper equation in point-slope form for the given point and slope. Point: (–3, 7); Slope: 4 (Points : 1) y + 7 = 4(x + 3) y + 7 = 4x – 12 y – 7 = 4(x – 3) y – 7 = 4(x + 3)

OpenStudy (thefluffmuncher):

i'm spilt between c and d

OpenStudy (danjs):

point (x1,y1) slope m y - y1 = m*(x-x1)

OpenStudy (danjs):

minus a negative is the same as addition

OpenStudy (thefluffmuncher):

so its d

OpenStudy (danjs):

(x - (-3)) = x +3

OpenStudy (danjs):

right

OpenStudy (thefluffmuncher):

thanks